Both the StanCam and FIES Cryotigers detector cooling systems have caused further problems over the past six months. In early February 2010 the StanCam detector temperature rose to -58 C (unfortunately there is no dewar temperature sensor for StanCam). No obvious problem with the cooling system could be found but the dewar vacuum was low, but this could have been due to the instrument warming up. The dewar was pumped and a bit more PT13 gas was put in the cooling system since its pressure was slightly below the preferred level. A couple of days after the temperature rose again, and again the dewar was pumped. Since the second pumping no large rises of the detector temperature have occurred.

The filter/dryer of the FIES system was replaced with a new one last November. Though it did seem to have helped at the time, this does not appear to avoid the occasional problems as described below.

As a result of the lightning storm in February the FIES dewar warmed up. After pumping the dewar and restarting the CryoTiger the cold-head temperature only fell to -160C and at this level it was not possible to maintain the reference temperature of -120 C for the detector. As a temporary measure the detector reference temperature was increased from -120 to -110 C. To try and recover the correct operating temperature the dewar was allowed to warm-up to room temperature pumped for a day and cooled but it only returned to -160C. On 30 March for no obvious reason the cold-head temperature dropped on its own to -199C, then a couple of days later rose to -196C and has been stable since. The detector reference temperature was put back to -120C.

Thomas Augusteijn 2010-05-27